Heavy cream, which is 30% butterfat, and lowfat milk, which is 2% butterfat, are added together to make 100 gallons of regular milk, which is 4% butterfat. How much of cream and low fat milk should be used?
This is what I have set up (i don't know if it's right...):
.30x + .02y = .04(100)

Answer by rapaljer(4671)

This is correct, except that if you have two variables you will need two equations. The other equation (the easy one!) was x+ y = 100, from which you might want to solve for y = 100-x, and substitute into the other equation to get:

.30x + .02(100-x) = .04(100)
.30x + 2 - .02x = 4
.28x = 2
x = 2/.28 = 50/7 or approximately 7.14 gallons of cream
100-x = 92 6/7 or approximately 92.86 gallons of 2%

R^2 at SCC